# DexGrasp-Diffusion: Diffusion-based Unified Functional Grasp Synthesis Pipeline for Multi-Dexterous Robotic Hands

### v1

@@ Abstract
The versatility and adaptability of human grasping catalyze advancing dexterous robotic manipulation. While significant strides have been made in dexterous grasp generation, current research endeavors pivot towards optimizing object manipulation while ensuring functional integrity, emphasizing the synthesis of functional grasps following desired affordance instructions. This paper addresses the challenge of synthesizing functional grasps tailored to diverse dexterous robotic hands by proposing DexGrasp-Diffusion, an end-to-end modularized diffusion-based pipeline. DexGrasp-Diffusion integrates MultiHandDiffuser, a novel unified data-driven diffusion model for multi-dexterous hands grasp estimation, with DexDiscriminator, which employs a Physics Discriminator and a Functional Discriminator with open-vocabulary setting to filter physically plausible functional grasps based on object affordances. The experimental evaluation conducted on the MultiDex dataset provides substantiating evidence supporting the superior performance of MultiHandDiffuser over the baseline model in terms of success rate, grasp diversity, and collision depth. Moreover, we demonstrate the capacity of DexGrasp-Diffusion to reliably generate functional grasps for household objects aligned with specific affordance instructions.
I Introduction
@@
Figure 1: Overview of the diffusion sampling process with fixed target objects for multi-dexterous robotic hands performed by our presented DexGrasp-Diffusion pipeline. (a) EZGripper with apple. (b) Barrett with power drill. (c) Robotiq-3F with camera. (d) Allegro with water bottle. (e) ShadowHand with wine glass.
@@
In response to the aforementioned challenges, we propose DexGrasp-Diffusion, an end-to-end modularized functional grasp synthesis pipeline that combines multi-dexterous hand grasp estimation with open-vocabulary affordance detection to enhance the adaptability and manipulation abilities of dexterous hands for complex tasks (Fig. 1). DexGrasp-Diffusion includes MultiHandDiffuser, a novel unified data-driven diffusion model that samples multi-dexterous hand grasps, and DexDiscriminator, which consists of a Physics Discriminator and a Functional Discriminator. Given an object 3D point cloud, our MultiHandDiffuser first generates diverse robust grasp poses of one specific hand. Thereafter, DexDiscriminator will eliminate physically invalid candidates and select suitable and feasible functional grasps associated with the desired affordances. We conduct experiments to assess DexGrasp-Diffusion’s ability to generate physically plausible functional grasps on the MultiDex dataset [1], which encompasses a varied array of grasp poses tailored for multiple dexterous robotic hands ranging from two to five fingers.
@@
    We integrate MultiHandDiffuser with DexDiscriminator and propose an effective modularized pipeline, DexGrasp-Diffusion, for generating feasible and reasonable functional grasps based on desired object affordances with open-vocabulary setting.
II Related Works
II-A Data-Driven Dexterous Grasping
@@
II-B Diffusion Models for Robotics
@@
Figure 2: Overview of our proposed DexGrasp-Diffusion pipeline.
III Method
III-A Problem Statement
In this work, we assume that the representation of the given object o $\in\mathbb{R}^{N\times 3}$ is point cloud. Let h = (t, $\theta$) represents a posture of the hand, where $\textbf{t}\in{\mathbb{R}}^{3}$ denotes global translation, $\theta\in\mathbb{R}^{k}$ represents the joint angles of the hand model, with $k$ denoting the degrees of freedom (DoF). Object point cloud o is first randomly rotated within its own frame, subsequently translation t and joint angles $\theta$ are denoised with MultiHandDiffuser (Sec. III-B) to match the robotic hand with the rotated object point cloud, generating a diverse set =$\mathbb{H}={\{\textbf{h}_{i}}\}_{i=1}^{m}$ of hand postures within object frame. Afterwards, the generated hand postures are filtered by Physics Discriminator (Sec. III-C) for physically stable grasping postures and Functional Discriminator (Sec. III-D) for suitable functional grasps.
III-B MultiHandDiffuser
@@ III-D Functional Discriminator
Table I: A comparison between different models on the MultiDex dataset. Succ.: Success Rate (%$\uparrow$. Div.: Diversity (rad.$\uparrow$. Col.: Collision Depth (mm)$\downarrow$. $\uparrow$: The higher, the better. $\downarrow$: The lower, the better. : Indicates the second-best result.
@@
IV Experiments
In this section, we undertake a series of experiments aimed at elucidating the efficacy of our proposed DexGrasp-Diffusion pipeline on the MultiDex dataset. Initially, we commence by comparing our approach with SceneDiffuser baseline models, which are exclusively trained using single-hand data. Subsequently, we furnish diverse ablation studies to facilitate a comprehensive investigation of the MultiHandDiffuser. Thirdly, we showcase noteworthy qualitative outcomes attained through DexGrasp-Diffusion. Lastly, an analysis of failure instances and prospective research directions is deliberated.
IV-A Dataset
IV-B Evaluation Metrics
IV-C Implementation Details
V Results and Analysis
V-A Quantitative Analysis
Baselines. We adopt SceneDiffuser as our baseline model. Since it is designed for grasp generating with single-hand model, we trained SceneDiffuser on each subset of the MultiDex dataset and obtained five distinct models to handle each hand model. Our proposed MultiHandDiffuser is directly trained on the Multidex dataset for all five dexterous hands. The evaluation result, as shown in Tab. I, demonstrates that our model achieves higher mean accuracy, higher grasp diversity, and less collision compared to baseline models. This superior performance not only ensures better overall efficacy but also underscores the model’s enhanced capability to generalize across multi-dexterous hands. It is noteworthy that the CAD model of the Robotiq-3F from the MultiDex dataset experiences severe self-collisions within the Isaac Gym environment, which cause inaccurate evaluation data with the Physics Discriminator as these collisions displace objects. Thus, results related to the Robotiq-3F are excluded from our analysis.
@@ V-A Quantitative Analysis
The second set is denoted as pc, which means we discard hand class and finger label and only extract hand embedding on unlabeled hand point cloud. Performance on both success rate and collision drops as the network is no longer explicitly informed which type of hand it is dealing with, making it more challenging for both training and inference. The third set is denoted as pc+handclass, which represents using unlabeled hand point cloud as condition while informing the network which hand it is handling. Results in Tab. I show that it not only outperforms pc but also outperforms handclass, meaning that the combination of hand point cloud and hand class enables the network to take advantage of diverse training samples while capturing the geometric features of the hand for better matching between objects and hands.
In fullset, all of the conditions including finger label are provided to the network, further enabling it to differentiate between multiple hands and different fingers. Therefore, fullset achieves the highest mean success rate, second-best and comparable diversity and collision depth. Furthermore, we systematically vary the diffusion time steps $T$ for fullset model and report the mean value of each metric in Tab. II. We observe that $T$ plays a crucial role in balancing the diversity and success rate of dexterous grasp estimation. Specifically, a value of $T$ = 100 yields optimal diversity in produced grasps, while $T$ = 1000 results in the highest mean success rate.
Table II: A comparison between different fullest models on the MultiDex dataset. Succ.: Success Rate (%). Div.: Diversity (rad.). Col.: Collision Depth (mm). $\uparrow$: The higher, the better. $\downarrow$: The lower, the better.
@@ V-B Qualitative Results
Generated Functional Grasps. In Fig. 4, we present several illustrative instances utilizing objects sourced from the MultiDex dataset, which effectively showcase DexGrasp-Diffusion’s capacity to generate functional grasps aligned with desired affordance labels. Notably, for simple and seen affordances in the training set of OpenAD [8], the attainment of high-quality affordance detection outcomes substantially facilitates the selection of dependable functional grasps by DexGrasp-Diffusion. Conversely, when confronted with challenging, previously unseen affordance labels, although the derived affordance regions may not exhibit absolute precision, our DexGrasp-Diffusion algorithm adeptly discerns and filters rational and suitable functional grasps, which proves the robustness of the proposed pipeline and its ability to handle different desired affordance labels of varying complexity.
V-C Discussion
Figure 5: Some failure or counter-intuitive cases of our pipeline. The unseen affordances are shown in orange.
Despite yielding promising results, it is imperative to acknowledge that our pipeline has not achieved flawless proficiency in multi-dexterous hand grasp synthesis and universal functional grasp detection. Instances, where DexGrasp-Diffusion manifests its limitations, are delineated in Fig. 5. Specifically, the first two cases depict instances wherein our MultiHandDiffuser failed to generate viable grasps for the bowl and pan, despite attempts by the hands to access and grasp the bottom regions of said objects. We believe that these failure examples stem from the absence of explicit collision constraints during the training phase of MultiHandDiffuser, leading to instances where the hand disregards potential collisions and proceeds to grasp the wrong positions. Furthermore, the scarcity of objects with similar shapes within the training dataset contributes to diminished success rates in MultiHandDiffuser’s grasp predictions for objects such as bowls and pans. Consequently, having a comprehensive largescale multi-dexterous hand dataset, encompassing objects with more intricate geometries would facilitate enhanced model training. Subsequently, an interesting yet counter-intuitive grasp is observed in the third case depicted in Fig. 5, wherein the model endeavors to grasp an apple with two fingers while extending another finger towards the opposing end. We attribute this phenomenon to analogous physically feasible but unreasonable ground truth grasp poses within the MultiDex dataset, thereby resulting in the learning of corresponding data distributions during the training of MultiHandDiffuser.
The final case exemplifies an occurrence where our pipeline generates a false-positive grasp, which fails to be a desired functional grasp according to the given affordance label due to the inaccurate and noisy affordance detection by Functional Discriminator. Furthermore, we also note that only the orientation of the input object point cloud is consistent with the orientation of the object in the OpenAD’s training set, the OpenAD model may obtain appropriate affordance detection results, which weakens the robustness of our pipeline to a certain extent. Given the modular nature of our pipeline, substituting OpenAD with a more adaptable and stable open-vocabulary 3D point cloud affordance detection module could ameliorate this limitation.
VI Conclusions

### v2

# Zhengshen Zhang, Lei Zhou, Chenchen Liu, Zhiyang Liu, Chengran Yuan, Sheng Guo, Ruiteng Zhao, Marcelo H. Ang Jr., and Francis EH Tay

@@ Abstract
The versatility and adaptability of human grasping catalyze advancing dexterous robotic manipulation. While significant strides have been made in dexterous grasp generation, current research endeavors pivot towards optimizing object manipulation while ensuring functional integrity, emphasizing the synthesis of functional grasps following desired affordance instructions. This paper addresses the challenge of synthesizing functional grasps tailored to diverse dexterous robotic hands by proposing DexGrasp-Diffusion, an end-to-end modularized diffusion-based method. DexGrasp-Diffusion integrates MultiHandDiffuser, a novel unified data-driven diffusion model for multi-dexterous hands grasp estimation, with DexDiscriminator, which employs a Physics Discriminator and a Functional Discriminator with open-vocabulary setting to filter physically plausible functional grasps based on object affordances. The experimental evaluation conducted on the MultiDex dataset provides substantiating evidence supporting the superior performance of MultiHandDiffuser over the baseline model in terms of success rate, grasp diversity, and collision depth. Moreover, we demonstrate the capacity of DexGrasp-Diffusion to reliably generate functional grasps for household objects aligned with specific affordance instructions.
keywords:
dexterous grasping, affordance detection, diffusion model
1 Introduction
@@
Figure 1: Overview of the diffusion sampling process with fixed target objects for multi-dexterous robotic hands performed by our presented DexGrasp-Diffusion method. (a) EZGripper with apple. (b) Barrett with power drill. (c) Robotiq-3F with camera. (d) Allegro with water bottle. (e) ShadowHand with wine glass.
@@
In response to the aforementioned challenges, we propose DexGrasp-Diffusion, an end-to-end modularized functional grasp synthesis method that combines multi-dexterous hand grasp estimation with open-vocabulary affordance detection to enhance the adaptability and manipulation abilities of dexterous hands for complex tasks (Fig. 1). DexGrasp-Diffusion includes MultiHandDiffuser, a novel unified data-driven diffusion model that samples multi-dexterous hand grasps, and DexDiscriminator, which consists of a Physics Discriminator and a Functional Discriminator. Given an object 3D point cloud, our MultiHandDiffuser first generates diverse robust grasp poses of one specific hand. Thereafter, DexDiscriminator will eliminate physically invalid candidates and select suitable and feasible functional grasps associated with the desired affordances. We conduct experiments to assess DexGrasp-Diffusion’s ability to generate physically plausible functional grasps on the MultiDex dataset [1], which encompasses a varied array of grasp poses tailored for multiple dexterous robotic hands ranging from two to five fingers.
@@
    We integrate MultiHandDiffuser with DexDiscriminator and propose an effective modularized approach, DexGrasp-Diffusion, for generating feasible and reasonable functional grasps based on desired object affordances with open-vocabulary setting.
2 Related Works
2.1 Data-Driven Dexterous Grasping
@@
2.2 Diffusion Models for Robotics
@@
Figure 2: Overview of our proposed DexGrasp-Diffusion method.
3 Method
3.1 Problem Statement
In this work, we assume that the representation of the given object o $\in\mathbb{R}^{N\times 3}$ is point cloud. Let h = (t, $\theta$) represents a posture of the hand, where $\textbf{t}\in{\mathbb{R}}^{3}$ denotes global translation, $\theta\in\mathbb{R}^{k}$ represents the joint angles of the hand model, with $k$ denoting the degrees of freedom (DoF). Object point cloud o is first randomly rotated within its own frame, subsequently translation t and joint angles $\theta$ are denoised with MultiHandDiffuser (Sec. 3.2) to match the robotic hand with the rotated object point cloud, generating a diverse set =$\mathbb{H}={\{\textbf{h}_{i}}\}_{i=1}^{m}$ of hand postures within object frame. Afterwards, the generated hand postures are filtered by Physics Discriminator (Sec. 3.3) for physically stable grasping postures and Functional Discriminator (Sec. 3.4) for suitable functional grasps.
3.2 MultiHandDiffuser
@@ 3.4 Functional Discriminator
4 Experiments
In this section, we undertake a series of experiments aimed at elucidating the efficacy of our proposed DexGrasp-Diffusion approach on the MultiDex dataset. Initially, we commence by comparing our approach with SceneDiffuser baseline models, which are exclusively trained using single-hand data. Subsequently, we furnish diverse ablation studies to facilitate a comprehensive investigation of the MultiHandDiffuser. Thirdly, we showcase noteworthy qualitative outcomes attained through DexGrasp-Diffusion. Lastly, an analysis of failure instances and prospective research directions is deliberated.
4.1 Dataset
4.2 Evaluation Metrics
4.3 Implementation Details
5 Results and Analysis
5.1 Quantitative Analysis
Baselines. We adopt SceneDiffuser as our baseline model. Since it is designed for grasp generating with single-hand model, we trained SceneDiffuser on each subset of the MultiDex dataset and obtained five distinct models to handle each hand model. Our proposed MultiHandDiffuser is directly trained on the Multidex dataset for all five dexterous hands. The evaluation result, as shown in Tab. 1, demonstrates that our model achieves higher mean accuracy, higher grasp diversity, and less collision compared to baseline models. This superior performance not only ensures better overall efficacy but also underscores the model’s enhanced capability to generalize across multi-dexterous hands. It is noteworthy that the CAD model of the Robotiq-3F from the MultiDex dataset experiences severe self-collisions within the Isaac Gym environment, which cause inaccurate evaluation data with the Physics Discriminator as these collisions displace objects. Thus, results related to the Robotiq-3F are excluded from our analysis.
Table 1: A comparison between different models on the MultiDex dataset. Succ.: Success Rate (%$\uparrow$. Div.: Diversity (rad.$\uparrow$. Col.: Collision Depth (mm)$\downarrow$. $\uparrow$: The higher, the better. $\downarrow$: The lower, the better. : Indicates the second-best result.
@@ 5.1 Quantitative Analysis
The second set is denoted as pc, which means we discard hand class and finger label and only extract hand embedding on unlabeled hand point cloud. Performance on both success rate and collision drops as the network is no longer explicitly informed which type of hand it is dealing with, making it more challenging for both training and inference. The third set is denoted as pc+handclass, which represents using unlabeled hand point cloud as condition while informing the network which hand it is handling. Results in Tab. 1 show that it not only outperforms pc but also outperforms handclass, meaning that the combination of hand point cloud and hand class enables the network to take advantage of diverse training samples while capturing the geometric features of the hand for better matching between objects and hands.
In fullset, all of the conditions including finger label are provided to the network, further enabling it to differentiate between multiple hands and different fingers. Therefore, fullset achieves the highest mean success rate, second-best and comparable diversity and collision depth. Furthermore, we systematically vary the diffusion time steps $T$ for fullset model and report the mean value of each metric in Tab. 2. We observe that $T$ plays a crucial role in balancing the diversity and success rate of dexterous grasp estimation. Specifically, a value of $T$ = 100 yields optimal diversity in produced grasps, while $T$ = 1000 results in the highest mean success rate.
Table 2: A comparison between different fullest models on the MultiDex dataset. Succ.: Success Rate (%). Div.: Diversity (rad.). Col.: Collision Depth (mm). $\uparrow$: The higher, the better. $\downarrow$: The lower, the better.
@@ 5.2 Qualitative Results
Generated Functional Grasps. In Fig. 4, we present several illustrative instances utilizing objects sourced from the MultiDex dataset, which effectively showcase our pipeline’s capacity to generate functional grasps aligned with desired affordance labels. Notably, for simple and seen affordances in the training set of OpenAD [8], the attainment of high-quality affordance detection outcomes substantially facilitates the selection of dependable functional grasps by DexGrasp-Diffusion. Conversely, when confronted with challenging, previously unseen affordance labels, although the derived affordance regions may not exhibit absolute precision, our DexGrasp-Diffusion algorithm adeptly discerns and filters rational and suitable functional grasps, which proves the robustness of the proposed method and its ability to handle different desired affordance labels of varying complexity.
@@
5.3 Discussion
Figure 5: Some failure or counter-intuitive cases of our method. The unseen affordances are shown in orange.
Despite yielding promising results, it is imperative to acknowledge that our method has not achieved flawless proficiency in multi-dexterous hand grasp synthesis and universal functional grasp detection. Instances, where DexGrasp-Diffusion manifests its limitations, are delineated in Fig. 5. Specifically, the first two cases depict instances wherein our MultiHandDiffuser failed to generate viable grasps for the bowl and pan, despite attempts by the hands to access and grasp the bottom regions of said objects. We believe that these failure examples stem from the absence of explicit collision constraints during the training phase of MultiHandDiffuser, leading to instances where the hand disregards potential collisions and proceeds to grasp the wrong positions. Furthermore, the scarcity of objects with similar shapes within the training dataset contributes to diminished success rates in MultiHandDiffuser’s grasp predictions for objects such as bowls and pans. Consequently, having a comprehensive largescale multi-dexterous hand dataset, encompassing objects with more intricate geometries would facilitate enhanced model training. Subsequently, an interesting yet counter-intuitive grasp is observed in the third case depicted in Fig. 5, wherein the model endeavors to grasp an apple with two fingers while extending another finger towards the opposing end. We attribute this phenomenon to analogous physically feasible but unreasonable ground truth grasp poses within the MultiDex dataset, thereby resulting in the learning of corresponding data distributions during the training of MultiHandDiffuser.
The final case exemplifies an occurrence where our approach generates a false-positive grasp, which fails to be a desired functional grasp according to the given affordance label due to the inaccurate and noisy affordance detection by Functional Discriminator. Furthermore, we also note that only the orientation of the input object point cloud is consistent with the orientation of the object in the OpenAD’s training set, the OpenAD model may obtain appropriate affordance detection results, which weakens the robustness of our method to a certain extent. Given the modular nature of our method, substituting OpenAD with a more adaptable and stable open-vocabulary 3D point cloud affordance detection module could ameliorate this limitation.
6 Conclusions